A tobacco company produces blends of tobacco, with each blend containing various proportions of Turkish, domestic, and other tobaccos. The proportions of Turkish and domestic in a blend are random variables with joint density function (X = Turkish and Y = domestic) f(x,y)={(3x-y)/9, 1<x<3 , 1<y<2, 0, elsewhere a) Find the marginal density functions of X and Y. b) Find the mean and variance of X and Y c) Are X and Y independent?
1
Expert's answer
2020-05-11T18:36:51-0400

a)


"g(x)=\displaystyle\int_{-\infin}^{\infin}f(x,y)dy=\displaystyle\int_{1}^{2}{3x-y\over 9}dy="

"=\bigg[{6xy-y^2 \over 18}\bigg]\begin{matrix}
 2 \\
 1
\end{matrix}={12x-4-6x+1\over 18}={2x-1\over 6}"

"={x\over 3}-{1\over 6},\ for\ 1<x<3"

"h(y)=\displaystyle\int_{-\infin}^{\infin}f(x,y)dx=\displaystyle\int_{1}^{3}{3x-y\over 9}dx="

"=\bigg[{3x^2-2xy \over 18}\bigg]\begin{matrix}
 3 \\
 1
\end{matrix}={27-6y-3+2y\over 18}={12-2y\over 9}="

"={4\over 3}-{2y\over 9},\ for\ 1<y<2"

b)


"E(X)=\displaystyle\int_{-\infin}^{\infin}xg(x)dx=\displaystyle\int_{1}^{3}({x^2\over 3}-{x\over 6})dx="

"=\bigg[{x^3 \over 9}-{x^2 \over 12}\bigg]\begin{matrix}
 3 \\
 1
\end{matrix}={27 \over 9}-{9 \over 12}-({1 \over 9}-{1 \over 12})={20\over 9}"


"E(X^2)=\displaystyle\int_{-\infin}^{\infin}x^2g(x)dx=\displaystyle\int_{1}^{3}({x^3\over 3}-{x^2\over 6})dx="

"=\bigg[{x^4 \over 12}-{x^3 \over 18}\bigg]\begin{matrix}
 3 \\
 1
\end{matrix}={81 \over 12}-{27 \over 18}-({1 \over 12}-{1 \over 18})={47\over 9}"

"V(X)=E(X^2)-(E(X))^2={47\over 9}-({20\over 9})^2={23\over 81}"


"E(Y)=\displaystyle\int_{-\infin}^{\infin}yh(y)dy=\displaystyle\int_{1}^{2}({4y\over 3}-{2y^2\over 9})dy="

"=\bigg[{4y^2 \over 6}-{2y^3 \over 27}\bigg]\begin{matrix}
 2 \\
 1
\end{matrix}={16 \over 6}-{16 \over 27}-({4 \over 6}-{2 \over 27})={40\over 27}"

"E(Y^2)=\displaystyle\int_{-\infin}^{\infin}y^2h(y)dy=\displaystyle\int_{1}^{2}({4y^2\over 3}-{2y^3\over 9})dy="

"=\bigg[{4y^3 \over 9}-{2y^4 \over 36}\bigg]\begin{matrix}
 2 \\
 1
\end{matrix}={32 \over 9}-{32 \over 36}-({4 \over 9}-{2 \over 36})={41\over 18}"

"V(Y)=E(Y^2)-(E(Y))^2={41\over 18}-({40\over 27})^2={121\over 1458}"

c) The random variables X and Y are said to be statistically independent if and only if


"f(x, y)=g(x)h(y)"

for all "(x, y)" within their range.


"g(x)h(y)={2x-1\over 6}\cdot{12-2y\over 9}="

"={24x-4xy-12+2y\over 54}={12x-2xy-6+y\over 27}\not={3x-y\over 9}=f(x,y)"

Let "x=1.5, y=1.25"


"g(1.5)={2(1.5)-1\over 6}={1\over 3}"

"h(1.25)={12-2(1.25)\over 9}={19\over 18}"

"g(1.5)h(1.25)={1\over 3}({19\over 18})={19\over 54}"

"f(1.5,1.25)={3(1.5)-1.25\over 9}={13\over 36}\not={19\over 54}"

So the random variables "X" and "Y" are not independent.